OpenStudy (anonymous):

solve each system by substitution be sure to give your solution as an ordered pair 7x+3y=35 x-2y=5

OpenStudy (anonymous):

Can you multiply the second equation by -7 and then sum them?

OpenStudy (anonymous):

i tried

OpenStudy (anonymous):

i ended up with a decimal when i got to the end

OpenStudy (anonymous):

i got -2.714 etc etc

OpenStudy (bahrom7893):

7x+3y=35 x=5+2y 35+14y+3y=35 17y=0 y=0 7x+0=35 x=5

OpenStudy (anonymous):

oooh... i messed up big time.. thank you so much!